The difference between a number and 21 1/2 is no more than 14 1/4
Answer by Alan3354(69443)

The difference between a number and 21 1/2 is no more than 14 1/4
---------------
The difference between a number and 21.5 is no more than 14.25
--
n - 21.5 <= 14.25
Or
21.5 - n <= 14.25
------------------------------
--> |n - 21.5| <= 14.25 --- absolute value